OpenStudy (anonymous):

solve the equation (x-9)(4x+5)=0

OpenStudy (anonymous):

What do you think it could be...?

OpenStudy (anonymous):

\[x = -5/4\]

OpenStudy (anonymous):

or x= 9

OpenStudy (anonymous):

is it right

OpenStudy (anonymous):

You have to explain how u got your answers:-)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

becuse (x-9)*(4x +5) = 0 then either x- 9 = 0 or 4x+5=0 so x= 9 or x = -5/4

OpenStudy (bahrom7893):

Okay the only way a product can be 0 is if any of the terms are 0.. oh well problemsolver got there first..

OpenStudy (anonymous):

X = -5/4 and x = 9 i got this answer becauseif you want the equation to be equal then one of the terms must equal zero, you take the opposite. (x-9) so x=9. and (4x+5) so it would be -5/4

OpenStudy (anonymous):

thank u...

OpenStudy (anonymous):

yep